OpenStudy (anonymous):

Tracy needs to find the length and width of the equipment storage room where she works. the room is 2 meters longer than it is wide. Its perimeter is 16 meters. Find the width in meters.

OpenStudy (anonymous):

L=7

OpenStudy (anonymous):

Perimeter is the sum of all sides. A rectangle has 4 sides, 2 width and 2 length. So, P=24=2W+2L We also know that the length and width can be related by L=W+2 Replace L with W+2 in the original equation. P=24=2W+2L=2W+2(W+2)=4W+4 Solve for W 20=4W W=5 L=7

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

I do. Thank you so much! I've never been good at perimeter problems.

OpenStudy (anonymous):

Yeah, thats how I was at first. Your welcome.

OpenStudy (anonymous):

Could you give me best response?

OpenStudy (anonymous):

Absolutely.

OpenStudy (anonymous):

Thanks